OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4.

OpenStudy (anonymous):

-2x-6=-2x-2-4 -2x-6=-2x-6 0=0 so x=0

OpenStudy (ahsome):

\[-2(x+3)=-2(x+1)-4\] Expand the brackets \[-2x-6=-2x-2-4\] Move the \(x\) value to the LHS \[-2x-6+2x=-2-4\] Simplify \[0-6=-2-4\] \[-6=-6\] Therefore, \(x\) equals 0

OpenStudy (anonymous):

Here are the options @Ahsome 2 3 all real number or no solution